The perimeter of a rectangle is 98 m. If you add 3 m to its length and 8 m to its width, you obtain a square. What are the original dimensions of this lot?
Answer by TimothyLamb(4379)

2x + 2y = 98
x + 3 = y + 8
---
2x + 2y = 98
x - y = 5
---
copy and paste the above standard form equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 27 m
y= 22 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php